# A discussion on the symmetry of relativistic Vlasov gas and its accretion in Kerr-Newman black hole

### v1

@@ IV.1 Observable Quantities
The particle flux density $J_{\mu}$ and the energy-momentum tensor $T_{\mu\nu}$ for Vlasov gases are defined by Eq. (4). We first explicitly formulate the phase space integration. The volume element
@@
utilizing $L\mathrm{d}L=\bar{L}\mathrm{d}\bar{L}$ and $\bar{f}_{n}\equiv\int^{+\infty}_{0}m^{n}f\mathrm{d}m$, where $p_{\mu}\leftarrow\frac{p_{\mu}}{m},E\leftarrow\frac{E}{m}$ and $\bar{L}\leftarrow\frac{\bar{L}}{m}$, hereafter denote quantities per unit mass. The momentum components are
@@
The integrals over absorption and scattering domains are
@@
$n_{J}:=\sqrt{-g^{ab}J_{a}J_{b}}.$ (36)
All subsequent analysis adopts the Locally Non-Rotating Frame (LNRF) with tetrad Bardeen(1972) ,
@@
$\varepsilon:=T_{ab}(e_{0})^{a}(e_{0})^{b}=e^{-2\nu}(T_{tt}+2\omega T_{t\varphi}+\omega^{2}T_{\varphi\varphi}).$ (39)
@@
The divergence-free energy-momentum tensor $\nabla^{a}\mathcal{T}_{ab}=0$ permits the construction of conserved currents $\varXi_{a}:=-\mathcal{T}_{ab}\xi^{b},\varPsi_{a}:=\mathcal{T}_{ab}\psi^{b}$. The mass, energy, and angular momentum accretion rates of the black hole are respectively defined as
$\dot{\mathscr{M}}$ $:=\int_{S}J_{a}\bar{n}^{a},$$\dot{\mathscr{E}}$ $:=\int_{S}\varXi_{a}\bar{n}^{a},$$\dot{\mathscr{L}}$ $:=\int_{S}\varPsi_{a}\bar{n}^{a},$ (42)
@@ IV.2 Asymptotic Analysis
$[T^{\pm}]^{\mu\ \mathrm{sct}}_{\ \nu}/[\alpha_{0}m^{5}]$ $=2\pi\cdot\operatorname{diag}\left[-\left(\frac{K_{1}(z)}{z}+\frac{3K_{2}(z)}{z^{2}}\right),\frac{K_{2}(z)}{z^{2}},\frac{K_{2}(z)}{z^{2}},\frac{K_{2}(z)}{z^{2}}\right],$ (46b)
where $K_{n}(z)$ denotes the modified Bessel functions of the second kind. All the other components asymptotically vanish. The asymptotic results of some physical quantities are
@@
$\varepsilon_{\infty}$ $=4\pi\left(\frac{K_{1}(z)}{z}+\frac{3K_{2}(z)}{z^{2}}\right),$ $\mathcal{P}_{i}$ $=4\pi\frac{K_{2}(z)}{z^{2}},\quad\text{for}\ i=1,2,3.$ (47b)
@@
The physical quantities involved in divergence and their corrections are as follows:
@@
where the prefactor $S_{\mathcal{B}}\equiv 4\pi(r^{2}_{H}+a^{2})$ represents the area of the event horizon $\mathcal{B}$. Considering the associated divergent factor $\frac{1}{\sqrt{\varDelta}}$ ( $e^{\mu_{1}}=\sqrt{\rho^{2}/\varDelta}$), we introduce auxiliary parameters to analyze the accretion rate $\dot{\mathcal{A}}(a,Q)$ more effectively:
@@ V Numerical Results
This section presents numerical implementations of the physical quantities discussed in Sec. III. For gas properties: $M=1$, $a=Q=0.1$, $\kappa=0,\pm 0.3$, $z=1$. For accretion rates: $a,Q\in[0,1]$. The Vlasov gas and black hole form a coupled system; $Q$ may vary but $\kappa$ is fixed.
@@
Figure 2: (a) The particle number density $n_{J}/n_{J\infty}$ of the gas’s comoving frame and (b) the particle number density density $n/n_{\infty}$ of the LNRF on the equatorial plane. Dashed lines in (b) show rescaled $\tilde{n}$.
@@
Plasma observables closely match those of neutral gas. Thus, the subsequent plots label only the plasma curves; the neutral gas profiles are shown with black dotted lines.
@@
Figure 3: (a) The energy density $\varepsilon$ and the rescaled energy $\tilde{\varepsilon}\equiv e^{-2\nu}\varepsilon$. (b) The principal pressures $\mathcal{P}_{i}$ and the rescaled radial pressure $\tilde{\mathcal{P}}_{1}\equiv e^{-2\mu_{1}}\mathcal{P}_{1}$.
@@
The following employs numerical methods to investigate the accretion rates of Kerr black holes, RN black holes, and extreme KN ($a^{2}+Q^{2}\to 1$) black holes, focusing on their dependence on the rotation parameter $a$ and electric charge $Q$. Specific accretion rates are denoted as
$\dot{\mathcal{A}}_{S}$ $\equiv\dot{\mathcal{A}}(0,0),$ Kerr $\equiv\dot{\mathcal{A}}(a,0),$ $\text{KN}_{a}$ $\equiv\dot{\mathcal{A}}\left(a,\sqrt{1-a^{2}}\right),$
RN $\equiv\dot{\mathcal{A}}(0,1-Q),$ $\text{KN}_{Q}$ $\equiv\dot{\mathcal{A}}\left(\sqrt{1-Q^{2}},1-Q\right),$ (51)
@@
Figure 4: The mass accretion rate parameters of characteristic black holes relative to the Schwarzschild case are presented.
@@
Figure 5: The energy accretion rate parameters of characteristic black holes relative to the Schwarzschild case are presented.
@@
Figs. 4a and 4b demonstrate distinct rotational dependencies of Kerr black hole accretion rates: the mass accretion rate ˙$\dot{\mathcal{M}}_{\varDelta}$ decreases with growing rotation parameter $a$, while the average accretion rate per unit area ˙$\dot{\mathcal{M}}_{\delta}$ increases. These results align with the theoretical frameworks in Cieslik-Mach(2022) and Li(2023) respectively. The mass accretion ˙$\dot{\mathcal{M}}_{\varDelta}$ of RN black holes decreases with increasing charge parameter $Q$. However, the accretion rate parameter ˙$\dot{\mathcal{M}}_{\delta}$ exhibits distinct behaviors: it follows a non-monotonic trend with $Q$—initially decreasing before rising over the full parameter range. Extreme Kerr-Newman black holes exhibit contrasting behavior — both accretion rate parameters increase monotonically with $a$ (equivalently decreasing $Q$).
@@
Figure 6: The angular momentum accretion rate parameters of characteristic black holes relative to the particle number density $n_{J\infty}$ are presented.
Fig. 6 shows angular momentum accretion vanishes for non-rotating black holes ($a=0$). For Kerr and extremal KN, the angular momentum accretion rate parameters ˙$-\dot{\mathcal{L}}_{\Delta}$ and ˙$-\dot{\mathcal{L}}_{\delta}$ increase monotonically with $a$. Unlike mass/energy rates, Kerr’s angular momentum curve lies inside extremal KN’s loop-shaped curves.
@@ VI Conclusion
In this paper, we have studied the properties and specific applications of the Vlasov gas distribution function in Kerr spacetime. The main conclusions are as follows.
@@
In the LNRF, we have computed the particle number density $n$, the energy density $\varepsilon$, the principal pressures $\mathcal{P}_{i}$, and the accretion rates, resolving horizon divergences via rescaling. For a plasma governed by the Jẗtner distribution, we have explicitly calculated the asymptotic expressions at spatial infinity. Numerically, all these quantities decrease monotonically with $r$, matching Schwarzschild asymptotics. At finite $r$, the plasma and the neutral gas observables nearly coincide; the positively (negatively) charged gas’s components are suppressed (enhanced).
These characteristic black holes reveal nearly identical curves between mass accretion rate parameters (˙$(\dot{\mathcal{M}}_{\varDelta},\dot{\mathcal{M}}_{\delta})$ and energy accretion rate parameters (˙$(\dot{\mathcal{E}}_{\varDelta},\dot{\mathcal{E}}_{\delta})$. An increase in both the rotation parameter $a$ and charge parameter $Q$ suppresses the mass accretion rate ˙$\dot{\mathcal{M}}_{\varDelta}$ (equivalently, the energy accretion rate ˙$\dot{\mathcal{E}}_{\varDelta}$). In contrast, the angular momentum accretion rate ˙$\dot{\mathcal{L}}_{\varDelta}$ exhibits different dependencies: it increases with rising $a$ but decreases with increasing $Q$.
Acknowledgments
This work was supported by the Young Natural Science Foundation of Wuhan Donghu College. I sincerely thank my colleague Yu-qin Yao for computational assistance in generating the numerical data for $J_{\mu}$ and $T_{\mu\nu}$ through execution of predefined computational workflows.
@@ Appendix A The Hamiltonian Vector Fields
For the Killing vector field $K^{a}$ on $M$ with its induced one-parameter isometry group $\varPhi\equiv\left\{\phi_{k}\right\}$, the identity $\mathscr{L}_{K}K=[K,K]=0$ yields $\phi_{*}[K|_{x}]=K|_{\phi(x)}$. When $x$ lies on a particle trajectory $C(\tau)$, $\phi_{k}(x)$ resides on trajectory $C(k,\tau)$. This implies
@@
The symplectic form $\varOmega_{AB}$, defined as the exterior derivative of the Poincaré 1-form $\Theta_{A}\equiv p_{\mu}(\mathrm{d}x^{\mu})_{A}$, satisfies $\varOmega_{AB}=(\mathrm{d}p_{\mu})_{A}\wedge(\mathrm{d}x^{\mu})_{B}$. For charged particles with canonical momentum $\pi_{a}\equiv\pi_{\mu}(\mathrm{d}x^{\mu})_{a}$, the symplectic form in conjugate coordinates $\left\{x^{\mu},\pi_{\nu}\right\}$ becomes (seeAcuna(2022))
@@
where $\hat{F}_{AB}=\frac{1}{2}F_{\mu\nu}(\mathrm{d}x^{\mu})_{A}\wedge(\mathrm{d}x^{\nu})_{B}$. The Hamiltonian vector field $Z_{F}^{A}$ for a function $\mathcal{F}$ satisfies
$\mathrm{d}\mathcal{F}=-\mathrm{i}_{Z_{H}}\varOmega,$ (59)
@@
$\mathrm{i}_{Z}\varOmega:=Z^{A}\varOmega_{AB}=-X^{\alpha}(\mathrm{d}p_{\alpha})_{B}+(Y_{\beta}+qX^{\gamma}F_{\gamma\beta})(\mathrm{d}x^{\beta})_{A}.$ (60)
The Hamiltonian vector fields for these functions $H=\frac{1}{2}g^{\alpha\beta}p_{\alpha}p_{\beta}$ and $\pi(K)\equiv K^{\gamma}\pi_{\gamma}$ are respectively
@@ Appendix B Periastron Analysis
For $\kappa=0$, the decoupled solutions of Eq. (28) become:
@@
where $B=A-2C^{2}\rho^{2}-2C\sqrt{1-A\rho^{2}+C^{2}\rho^{4}}$ with $A\equiv\frac{2}{\rho^{2}}-\frac{r-1}{r\varDelta}$ and $C\equiv-\frac{a\sin\theta\sin\sigma}{\rho^{2}\sqrt{\varDelta}}$. These parameters satisfy $A=B+2C\sqrt{1-\rho^{2}B}$. As $B\to 0$ (implying $E_{0}\to\infty,\bar{L}_{0}\to\infty$), we find $A=2C$, leading to the equation for $r_{\text{ph}}$:
@@
For $\kappa\neq 0$, decoupling involves solving quartics, but the $r_{\text{ph}}$ equation from $E\to\infty$, $\bar{\mathcal{L}}\to\infty$ matches the $\kappa=0$ case.
@@
reducing to $r_{\mathrm{ph}}=3$ for the Schwarzschild black hole. For the extreme KN black holes $(a^{2}+Q^{2}\to 1,r_{H}\to 1$):
@@
Equatorial bound orbits $p_{\theta}=0\Rightarrow\sigma=\pm\frac{\pi}{2}$ yield prograde ($L_{z}>0$) and retrograde ($L_{z}<0$) photon radii:

### v2

@@ IV.1 Observable Quantities
The particle flux density $J_{\mu}$ and the energy-momentum tensor $T_{\mu\nu}$ for Vlasov gases are defined by Eq. (4). We first explicitly formulate the phase space integration. The volume element is
@@
utilizing $L\mathrm{d}L=\bar{L}\mathrm{d}\bar{L}$ and $\bar{f}_{n}\equiv\int^{+\infty}_{0}m^{n}f\mathrm{d}m$, where $p_{\mu}\leftarrow\frac{p_{\mu}}{m},E\leftarrow\frac{E}{m}$ and $\bar{L}\leftarrow\frac{\bar{L}}{m}$, hereafter, denote quantities per unit mass. The momentum components are
@@
The integrals over the absorption and scattering domains are
@@
$n_{J}\coloneqq\sqrt{-g^{ab}J_{a}J_{b}}.$ (36)
All subsequent analysis adopts the Locally Non-Rotating Frame (LNRF) with the tetrad Bardeen(1972) ,
@@
$\varepsilon\coloneqq T_{ab}(e_{0})^{a}(e_{0})^{b}=e^{-2\nu}(T_{tt}+2\omega T_{t\varphi}+\omega^{2}T_{\varphi\varphi}).$ (39)
@@
The divergence-free energy-momentum tensor $\nabla^{a}\mathcal{T}_{ab}=0$ permits the construction of conserved currents $\varXi_{a}\coloneqq-\mathcal{T}_{ab}\xi^{b},\varPsi_{a}\coloneqq\mathcal{T}_{ab}\psi^{b}$. The mass, energy, and angular momentum accretion rates of the black hole are defined as
$\dot{\mathscr{M}}$ $\coloneqq\int_{S}J_{a}\bar{n}^{a},$$\dot{\mathscr{E}}$ $\coloneqq\int_{S}\varXi_{a}\bar{n}^{a},$$\dot{\mathscr{L}}$ $\coloneqq\int_{S}\varPsi_{a}\bar{n}^{a},$ (42)
@@ IV.2 Asymptotic Analysis
$[T^{\pm}]^{\mu\ \mathrm{scat}}_{\ \nu}/[\alpha_{0}m^{5}]$ $=2\pi\cdot\operatorname{diag}\left[-\left(\frac{K_{1}(z)}{z}+\frac{3K_{2}(z)}{z^{2}}\right),\frac{K_{2}(z)}{z^{2}},\frac{K_{2}(z)}{z^{2}},\frac{K_{2}(z)}{z^{2}}\right],$ (46b)
where $K_{n}(z)$ denotes the modified Bessel functions of the second kind. All the other components asymptotically vanish. The asymptotic results for some physical quantities are
@@
$\varepsilon_{\infty}$ $=4\pi\left(\frac{K_{1}(z)}{z}+\frac{3K_{2}(z)}{z^{2}}\right),$ $\mathcal{P}_{i}$ $=4\pi\frac{K_{2}(z)}{z^{2}},\text{ for }i=1,2,3.$ (47b)
@@
The physical quantities exhibiting divergence and their corrections are as follows:
@@
where the prefactor $S_{\mathcal{B}}\equiv 4\pi(r^{2}_{H}+a^{2})$ represents the area of the event horizon $\mathcal{B}$. Considering the divergent factor $\frac{1}{\sqrt{\varDelta}}$ ( $e^{\mu_{1}}=\sqrt{\rho^{2}/\varDelta}$), we introduce auxiliary parameters to analyze the accretion rate $\dot{\mathcal{A}}(a,Q)$ more effectively:
@@ V Numerical Results
This section presents numerical implementations of the physical quantities discussed in Sec. III. For the gas properties: $M=1$, $a=Q=0.1$, $\kappa=0,\pm 0.3$, $z=1$. For the accretion rates: $a,Q\in[0,1]$. The Vlasov gas and the black hole form a coupled system; $Q$ may vary, but $\kappa$ is fixed.
@@
Figure 2: (a) The particle number density $n_{J}/n_{J\infty}$ of the comoving frame of the gas and (b) the particle number density $n/n_{\infty}$ in the LNRF on the equatorial plane. Dashed lines in (b) show the rescaled $\tilde{n}$.
@@
Plasma observables closely match those of neutral gas. Thus, in subsequent plots, only the plasma curves are labeled; the neutral gas profiles are shown with black dotted lines.
@@
Figure 3: (a) The energy density $\varepsilon$ and the rescaled energy density $\tilde{\varepsilon}\equiv e^{-2\nu}\varepsilon$. (b) The principal pressures $\mathcal{P}_{i}$ and the rescaled radial pressure $\tilde{\mathcal{P}}_{1}\equiv e^{-2\mu_{1}}\mathcal{P}_{1}$.
@@
The following employs numerical methods to investigate the accretion rates of Kerr black holes, RN black holes, and extreme KN ($a^{2}+Q^{2}\to 1$) black holes, focusing on their dependence on the rotation parameter $a$ and the electric charge $Q$. Specific accretion rates are denoted as
$\dot{\mathcal{A}}_{S}$ $\equiv\dot{\mathcal{A}}(0,0),$ $\mathrm{Kerr}$ $\equiv\dot{\mathcal{A}}(a,0),$ $\mathrm{KN}_{a}$ $\equiv\dot{\mathcal{A}}\left(a,\sqrt{1-a^{2}}\right),$
$\mathrm{RN}$ $\equiv\dot{\mathcal{A}}(0,1-Q),$ $\mathrm{KN}_{Q}$ $\equiv\dot{\mathcal{A}}\left(\sqrt{1-Q^{2}},1-Q\right),$ (51)
@@
Figure 4: Mass accretion rate parameters for characteristic black holes normalized by the Schwarzschild case are presented.
@@
Figure 5: Energy accretion rate parameters for characteristic black holes normalized by the Schwarzschild case are presented.
@@
Figs. 4a and 4b demonstrate distinct rotational dependencies of Kerr black hole accretion rates: the mass accretion rate ˙$\dot{\mathcal{M}}_{\varDelta}$ decreases with increasing rotation parameter $a$, while the average accretion rate per unit area ˙$\dot{\mathcal{M}}_{\delta}$ increases. These results align with the theoretical frameworks in Cieslik-Mach(2022) and Li(2023) , respectively. The mass accretion ˙$\dot{\mathcal{M}}_{\varDelta}$ of RN black holes decreases with increasing charge parameter $Q$. However, the accretion rate parameter ˙$\dot{\mathcal{M}}_{\delta}$ exhibits a non-monotonic trend with $Q$, initially decreasing before rising over the full parameter range. Extreme Kerr-Newman black holes exhibit contrasting behavior — both accretion rate parameters increase monotonically with $a$ (equivalently decreasing $Q$).
@@
Figure 6: Angular momentum accretion rate parameters for characteristic black holes normalized by the particle number density $n_{J\infty}$.
Fig. 6 shows that angular momentum accretion vanishes for non-rotating black holes ($a=0$). For Kerr and extremal KN black holes, the angular momentum accretion rate parameters ˙$-\dot{\mathcal{L}}_{\Delta}$ and ˙$-\dot{\mathcal{L}}_{\delta}$ increase monotonically with $a$. Unlike mass/energy rates, the angular momentum curve of Kerr black holes lies inside the loop-shaped curves of extremal KN black holes.
@@ VI Conclusion
In this paper, we have studied the properties and specific applications of the Vlasov gas distribution function in the KN spacetime. The main conclusions are as follows.
@@
In the LNRF, we have computed the particle number density $n$, the energy density $\varepsilon$, the principal pressures $\mathcal{P}_{i}$, and the accretion rates, resolving horizon divergences by rescaling. For a plasma governed by the Jüttner distribution, we have explicitly calculated the asymptotic expressions at spatial infinity. Numerically, all these quantities decrease monotonically with $r$, matching the Schwarzschild asymptotics. At finite $r$, the plasma and the neutral gas observables nearly coincide; components of the positively (negatively) charged gas are suppressed (enhanced).
For these characteristic black holes, we observe nearly identical curves between mass accretion rate parameters (˙$(\dot{\mathcal{M}}_{\varDelta},\dot{\mathcal{M}}_{\delta})$ and energy accretion rate parameters (˙$(\dot{\mathcal{E}}_{\varDelta},\dot{\mathcal{E}}_{\delta})$. An increase in both the rotation parameter $a$ and the charge parameter $Q$ suppresses the mass accretion rate ˙$\dot{\mathcal{M}}_{\varDelta}$ (equivalently, the energy accretion rate ˙$\dot{\mathcal{E}}_{\varDelta}$). In contrast, the angular momentum accretion rate ˙$-\dot{\mathcal{L}}_{\varDelta}$ exhibits different dependencies: it increases with rising $a$ but decreases with increasing $Q$.
Acknowledgements.
I sincerely thank my colleague Yu-qin Yao for computational assistance in generating the numerical data for $J_{\mu}$ and $T_{\mu\nu}$ by executing predefined computational workflows. Funding acknowledgment update:
The acknowledgment to the Young Natural Science Foundation of Wuhan Donghu College
in the original version (v1) has been removed due to force majeure.
@@ Appendix A The Hamiltonian Vector Fields
For a Killing vector field $K^{a}$ on $M$ with its induced one-parameter isometry group $\varPhi\equiv\left\{\phi_{k}\right\}$, the identity $\mathscr{L}_{K}K=[K,K]=0$ yields $\phi_{*}[K|_{x}]=K|_{\phi(x)}$. When $x$ lies on a particle trajectory $C(\tau)$, $\phi_{k}(x)$ resides on the trajectory $C(k,\tau)$. This implies
@@
The symplectic form $\varOmega_{AB}$, defined as the exterior derivative of the Poincaré 1-form $\Theta_{A}\equiv p_{\mu}(\mathrm{d}x^{\mu})_{A}$, satisfies $\varOmega_{AB}=(\mathrm{d}p_{\mu})_{A}\wedge(\mathrm{d}x^{\mu})_{B}$. For charged particles with canonical momentum $\pi_{a}\equiv\pi_{\mu}(\mathrm{d}x^{\mu})_{a}$, the symplectic form in conjugate coordinates $\left\{x^{\mu},\pi_{\nu}\right\}$ becomes (seeSarbach(2022) )
@@
where $\hat{F}_{AB}=\frac{1}{2}F_{\mu\nu}(\mathrm{d}x^{\mu})_{A}\wedge(\mathrm{d}x^{\nu})_{B}$. The Hamiltonian vector field $Z_{\mathcal{F}}^{A}$ of a function $\mathcal{F}$ satisfies
$\mathrm{d}\mathcal{F}=-\mathrm{i}_{Z_{\mathcal{F}}}\varOmega,$ (59)
@@
$\mathrm{i}_{Z}\varOmega\coloneqq Z^{A}\varOmega_{AB}=-X^{\alpha}(\mathrm{d}p_{\alpha})_{B}+(Y_{\beta}+qX^{\gamma}F_{\gamma\beta})(\mathrm{d}x^{\beta})_{A}.$ (60)
The Hamiltonian vector fields for the functions $H=\frac{1}{2}g^{\alpha\beta}p_{\alpha}p_{\beta}$ and $\pi(K)\equiv K^{\gamma}\pi_{\gamma}$ are respectively
@@ Appendix B Periastron Analysis
For $\kappa=0$, the decoupled solutions of Eq. (28) are
@@
where $B=A-2C^{2}\rho^{2}-2C\sqrt{1-A\rho^{2}+C^{2}\rho^{4}}$ with $A\equiv\frac{2}{\rho^{2}}-\frac{r-1}{r\varDelta}$ and $C\equiv-\frac{a\sin\theta\sin\sigma}{\rho^{2}\sqrt{\varDelta}}$. These parameters satisfy $A=B+2C\sqrt{1-\rho^{2}B}$. As $B\to 0$ (implying $E_{0}\to\infty,\bar{L}_{0}\to\infty$), we find $A=2C$, leading to the equation for $r_{\mathrm{ph}}$:
@@
For $\kappa\neq 0$, decoupling involves solving quartics, but the $r_{\mathrm{ph}}$ equation from $E\to\infty$, $\bar{\mathcal{L}}\to\infty$ matches the $\kappa=0$ case.
@@
reducing to $r_{\mathrm{ph}}=3$ for the Schwarzschild black hole. For extremal KN black holes $(a^{2}+Q^{2}\to 1,r_{H}\to 1$):
@@
Equatorial bound orbits ($p_{\theta}=0\Rightarrow\sigma=\pm\frac{\pi}{2}$) yield prograde ($L_{z}>0$) and retrograde ($L_{z}<0$) photon radii: